OpenStudy (goformit100):

A man takes a step forward with probability 0.4 and backward with probability 0.6. Find the probability that at the end of eleven steps he is one step away from from the starting point.

OpenStudy (anonymous):

question are getting briefer and briefer guess that is why they are called "lessens"

OpenStudy (goformit100):

@love_jessika15 @satellite73

OpenStudy (uri):

Combination..

OpenStudy (goformit100):

Ya we all here know we have to use it.... But How to use it @uri

OpenStudy (uri):

lol

OpenStudy (anonymous):

five forward and six back \[\binom{11}{5}.6^5.4^6\] i think

OpenStudy (uri):

One forword one backward So,11C6 or 11C5

OpenStudy (anonymous):

or five backward and six forward \[\binom{11}{5}.4^6.6^5\] add them up

OpenStudy (uri):

11C6=11*10*9*8*7*6/6*4*4*3*2*1= 11C5=11*10*9*8*7/5*4*3*2*1=

OpenStudy (goformit100):